# Modeling and Numerical Analysis of Kangaroo Lower Body based on Constrained Dynamics of Hybrid Serial-Parallel Floating-Base Systems

### 07 Dec 2023 v1

@@ A. Computation of Relative Kinematics
${{}^{a}\mathbf{\dot{R}}_{w}}$ $={{}^{a}\bm{\omega}_{w}}_{[\times]}{{}^{a}\mathbf{R}_{w}}=-{{}^{a}\bm{\omega}_{a}}_{[\times]}{{}^{a}\mathbf{R}_{w}}=$
$=-^{a}\mathbf{R}_{w}{{}^{w}\bm{\omega}_{a}}_{[\times]}{{}^{a}\mathbf{R}_{w}},$ (53)
@@
$=-{{}^{a}\mathbf{R}_{w}}{{}^{w}\bm{\omega}_{a}}\times{{}^{a}\mathbf{R}_{w}}{{}^{w}\bm{\omega}_{a,b}}+{{}^{a}\mathbf{R}_{w}}{{}^{w}\bm{\dot{\omega}}_{a,b}}=$
@@
Concerning the relative linear acceleration, its expression is reported in equation (56), obtained deriving equation (50a) w.r.t. time and using ${{}^{w}\mathbf{\hat{a}}_{a,b}}={{}^{w}\mathbf{a}_{b}}-{{}^{w}\mathbf{a}_{a}}$ and ${{}^{w}\mathbf{\hat{v}}_{a,b}}={{}^{w}\mathbf{v}_{b}}-{{}^{w}\mathbf{v}_{a}}$. Finally, from expression (56) we can compute the linear part of the relative Cartesian acceleration bias ${{}^{a}{\mathbf{\dot{J}}}_{a,b}^{\rightarrow}}\mathbf{\dot{q}}$, reported in equation (57) with $\mathbf{b}=-2{{}^{w}\bm{\omega}_{a}}\times{{}^{w}\mathbf{\hat{v}}_{a,b}}+{{}^{w}\bm{\omega}_{a}}\times{{}^{w}\bm{\omega}_{a}}\times{{}^{w}\mathbf{p}_{a,b}}$.
@@
${{}^{a}\mathbf{a}_{a,b}}$ $={{}^{a}\mathbf{\dot{R}}_{w}}{{}^{w}\mathbf{\hat{v}}_{a,b}}+{{}^{a}\mathbf{R}_{w}}{{}^{w}\mathbf{\hat{a}}_{a,b}}-{{}^{a}\mathbf{\dot{R}}_{w}}\left({{}^{w}\bm{\omega}_{a}}\times{{}^{w}\mathbf{p}_{a,b}}\right)-{{}^{a}\mathbf{R}_{w}}\left({{}^{w}\bm{\dot{\omega}}_{a}}\times{{}^{w}\mathbf{p}_{a,b}}\right)-{{}^{a}\mathbf{R}_{w}}\left({{}^{w}\bm{\omega}_{a}}\times{{}^{w}\mathbf{\hat{v}}_{a,b}}\right)=$
$={{}^{a}\mathbf{\dot{R}}_{w}}\left({{}^{w}\mathbf{\hat{v}}_{a,b}}-{{}^{w}\bm{\omega}_{a}}\times{{}^{w}\mathbf{p}_{a,b}}\right)+{{}^{a}\mathbf{R}_{w}}\left({{}^{w}\mathbf{\hat{a}}_{a,b}}-{{}^{w}\bm{\dot{\omega}}_{a}}\times{{}^{w}\mathbf{p}_{a,b}}-{{}^{w}\bm{\omega}_{a}}\times{{}^{w}\mathbf{\hat{v}}_{a,b}}\right)=$
$=-{{}^{a}\mathbf{R}_{w}}{{}^{w}\bm{\omega}_{a}}\times{{}^{a}\mathbf{R}_{w}}\left({{}^{w}\mathbf{\hat{v}}_{a,b}}-{{}^{w}\bm{\omega}_{a}}\times{{}^{w}\mathbf{p}_{a,b}}\right)+{{}^{a}\mathbf{R}_{w}}\left({{}^{w}\mathbf{\hat{a}}_{a,b}}-{{}^{w}\bm{\dot{\omega}}_{a}}\times{{}^{w}\mathbf{p}_{a,b}}-{{}^{w}\bm{\omega}_{a}}\times{{}^{w}\mathbf{\hat{v}}_{a,b}}\right)=$
$=-{{}^{a}\mathbf{R}_{w}}\left({{}^{w}\bm{\omega}_{a}}\times{{}^{w}\mathbf{\hat{v}}_{a,b}}-{{}^{w}\bm{\omega}_{a}}\times{{}^{w}\bm{\omega}_{a}}\times{{}^{w}\mathbf{p}_{a,b}}\right)+{{}^{a}\mathbf{R}_{w}}\left({{}^{w}\mathbf{\hat{a}}_{a,b}}-{{}^{w}\bm{\dot{\omega}}_{a}}\times{{}^{w}\mathbf{p}_{a,b}}-{{}^{w}\bm{\omega}_{a}}\times{{}^{w}\mathbf{\hat{v}}_{a,b}}\right)=$
$={{}^{a}\mathbf{R}_{w}}\left(-{{}^{w}\bm{\omega}_{a}}\times{{}^{w}\mathbf{\hat{v}}_{a,b}}+{{}^{w}\bm{\omega}_{a}}\times{{}^{w}\bm{\omega}_{a}}\times{{}^{w}\mathbf{p}_{a,b}}+{{}^{w}\mathbf{\hat{a}}_{a,b}}-{{}^{w}\bm{\dot{\omega}}_{a}}\times{{}^{w}\mathbf{p}_{a,b}}-{{}^{w}\bm{\omega}_{a}}\times{{}^{w}\mathbf{\hat{v}}_{a,b}}\right)=$
$={{}^{a}\mathbf{R}_{w}}\left({{}^{w}\mathbf{\hat{a}}_{a,b}}-{{}^{w}\bm{\dot{\omega}}_{a}}\times{{}^{w}\mathbf{p}_{a,b}}-2{{}^{w}\bm{\omega}_{a}}\times{{}^{w}\mathbf{\hat{v}}_{a,b}}+{{}^{w}\bm{\omega}_{a}}\times{{}^{w}\bm{\omega}_{a}}\times{{}^{w}\mathbf{p}_{a,b}}\right),$ (56)
@@
B. Projected Floating-Base Dynamics
@@
$\mathbf{M}_{b,b}(\mathbf{q})\begin{bmatrix}\mathbf{\ddot{p}}\\ \bm{\dot{\omega}}\end{bmatrix}+\mathbf{M}_{b,u}(\mathbf{q})\bm{\ddot{\theta}}_{u}+\mathbf{M}_{b,a}(\mathbf{q})\bm{\ddot{\theta}}_{a}+\mathbf{h}_{b}(\mathbf{q},\bm{\nu})$ $=\mathbf{J}_{c,b}(\mathbf{q})^{T}\mathbf{F}$ (60)
with $\mathbf{M}_{b,b}(\mathbf{q})\in\mathbb{R}^{6\times 6}$, $\mathbf{M}_{b,u}(\mathbf{q})\in\mathbb{R}^{6\times m}$ and $\mathbf{M}_{b,a}(\mathbf{q})\in\mathbb{R}^{6\times n-m}$. Substituting equation (3.1) and reorganizing leads to the floating-base part of the projected dynamics:
@@
Equations (61) and (64) permits to rewrite the QP problem in (3.2) only in the actuated and base accelerations, and contact forces.

### 22 Feb 2024 v2

@@ A-B Relative Acceleration and Acceleration Bias
${{}^{a}\mathbf{\dot{R}}_{w}}=-^{a}\mathbf{R}_{w}{{}^{w}\bm{\omega}_{a}}_{[\times]},$ (53)
@@
$=-{{}^{a}\mathbf{R}_{w}}{{}^{w}\bm{\omega}_{a}}\times{{}^{w}\bm{\omega}_{a,b}}+{{}^{a}\mathbf{R}_{w}}{{}^{w}\bm{\dot{\omega}}_{a,b}}=$
@@
Concerning the relative linear acceleration, its expression is reported in equation (56), obtained deriving equation (50a) w.r.t. time and using ${{}^{w}\mathbf{a}_{a,b}}={{}^{w}\mathbf{a}_{b}}-{{}^{w}\mathbf{a}_{a}}$ and ${{}^{w}\mathbf{v}_{a,b}}={{}^{w}\mathbf{v}_{b}}-{{}^{w}\mathbf{v}_{a}}$. Finally, from expression (56) we can compute the linear part of the relative Cartesian acceleration bias ${{}^{a}{\mathbf{\dot{J}}}_{a,b}^{\rightarrow}}\mathbf{\dot{q}}$, reported in equation (57) with $\mathbf{b}=-2{{}^{w}\bm{\omega}_{a}}\times{{}^{w}\mathbf{v}_{a,b}}+{{}^{w}\bm{\omega}_{a}}\times{{}^{w}\bm{\omega}_{a}}\times{{}^{w}\mathbf{p}_{a,b}}$.
@@
${{}^{a}\mathbf{a}_{a,b}}$ $={{}^{a}\mathbf{\dot{R}}_{w}}{{}^{w}\mathbf{v}_{a,b}}+{{}^{a}\mathbf{R}_{w}}{{}^{w}\mathbf{a}_{a,b}}-{{}^{a}\mathbf{\dot{R}}_{w}}\left({{}^{w}\bm{\omega}_{a}}\times{{}^{w}\mathbf{p}_{a,b}}\right)-{{}^{a}\mathbf{R}_{w}}\left({{}^{w}\bm{\dot{\omega}}_{a}}\times{{}^{w}\mathbf{p}_{a,b}}\right)-{{}^{a}\mathbf{R}_{w}}\left({{}^{w}\bm{\omega}_{a}}\times{{}^{w}\mathbf{v}_{a,b}}\right)=$
$=-{{}^{a}\mathbf{R}_{w}}{{}^{w}\bm{\omega}_{a}}\times{{}^{w}\mathbf{v}_{a,b}}+{{}^{a}\mathbf{R}_{w}}{{}^{w}\mathbf{a}_{a,b}}+{{}^{a}\mathbf{R}_{w}}{{}^{w}\bm{\omega}_{a}}\times\left({{}^{w}\bm{\omega}_{a}}\times{{}^{w}\mathbf{p}_{a,b}}\right)-{{}^{a}\mathbf{R}_{w}}\left({{}^{w}\bm{\dot{\omega}}_{a}}\times{{}^{w}\mathbf{p}_{a,b}}\right)-{{}^{a}\mathbf{R}_{w}}\left({{}^{w}\bm{\omega}_{a}}\times{{}^{w}\mathbf{v}_{a,b}}\right)=$
$={{}^{a}\mathbf{R}_{w}}\left[-{{}^{w}\bm{\omega}_{a}}\times{{}^{w}\mathbf{v}_{a,b}}+{{}^{w}\mathbf{a}_{a,b}}+{{}^{w}\bm{\omega}_{a}}\times\left({{}^{w}\bm{\omega}_{a}}\times{{}^{w}\mathbf{p}_{a,b}}\right)-\left({{}^{w}\bm{\dot{\omega}}_{a}}\times{{}^{w}\mathbf{p}_{a,b}}\right)-\left({{}^{w}\bm{\omega}_{a}}\times{{}^{w}\mathbf{v}_{a,b}}\right)\right]=$
$={{}^{a}\mathbf{R}_{w}}\left[{{}^{w}\mathbf{a}_{a,b}}-{{}^{w}\bm{\dot{\omega}}_{a}}\times{{}^{w}\mathbf{p}_{a,b}}-2{{}^{w}\bm{\omega}_{a}}\times{{}^{w}\mathbf{v}_{a,b}}+{{}^{w}\bm{\omega}_{a}}\times\left({{}^{w}\bm{\omega}_{a}}\times{{}^{w}\mathbf{p}_{a,b}}\right)\right],$ (56)
@@
Appendix B Projected Floating-Base Dynamics
@@
$\mathbf{M}_{b,b}(\mathbf{q})\begin{bmatrix}\mathbf{\ddot{p}}\\ \bm{\dot{\omega}}\end{bmatrix}+\mathbf{M}_{b,u}(\mathbf{q})\bm{\ddot{\theta}}_{u}+\mathbf{M}_{b,a}(\mathbf{q})\bm{\ddot{\theta}}_{a}+\mathbf{h}_{b}(\mathbf{q},\bm{\nu})$ $=$ (60)
$\mathbf{J}_{c,b}(\mathbf{q})^{T}\mathbf{F}$ (61)
with $\mathbf{M}_{b,b}(\mathbf{q})\in\mathbb{R}^{6\times 6}$, $\mathbf{M}_{b,u}(\mathbf{q})\in\mathbb{R}^{6\times m}$ and $\mathbf{M}_{b,a}(\mathbf{q})\in\mathbb{R}^{6\times n-m}$. Substituting equation (III-A) and reorganizing leads to the floating-base part of the projected dynamics:
@@
Equations (62) and (65) permits to rewrite the QP problem in (III-B) only in the actuated and base accelerations, and contact forces.